# Stochastic Processes and Transport Plans

### v2

@@ 1 Introduction
Generative models with neural network approximations have shown impressive results in many applications, in particular in inverse problems and data assimilation. First developments in this direction were generative adversarial networks by Goodfellow et al. [16] in 2014 and variational autoencoders by Kingma and Welling [24] in 2013. In this paper, we are interested in recent flow matching techniques pioneered by Liu et al. [28, 29] in 2022/23 from the point of view of stochastic processes, and by Lipman et al. [27] in 2023 via conditional probability paths. Flow matching is closely related to continuous normalizing flows (CNFs), also called neural ordinary differential equations by Chen et al. [9] in 2018, and to score-based diffusion models introduced by Sohl-Dickstein et al. [36] and Ermon and Song [38] in 2015 and 2019, respectively. The later ones can be considered as an instance of CNFs, but add the perspective of stochastic differential equations. To provide a more complete picture, we will briefly consider these approaches in Sections 10 and 11. Flow matching techniques stand out for their simple applicability and good scaling properties. They can be easily incorporated into other optimization techniques like plug-and-play approaches [30] and can be generalized for solving Bayesian inverse problems as we will see in Section 9.
@@
Our curves of interest are absolutely continuous curves in the Wasserstein space, which we introduce in Section 3. Due to the special metric
of the Wasserstein space, such a curve possesses a vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the curve-velocity pair $(\mu_{t},v_{t})$ fulfills the continuity equation
@@
    Here $v_{t}\mu_{t}=e_{t,\sharp}[(y-x)\alpha]$ is a vector valued measures, where you may note that for $\gamma\in\mathcal{P}_{2}(\mathbb{R}^{n}),f\in L^{1}(\mathbb{R}^{n},\gamma)$ the vector valued measure $f\gamma$ is defined by
    $\int_{\mathbb{R}^{n}}\varphi(x)\mathrm{d}f\gamma(x)=\int_{\mathbb{R}^{n}}\varphi(x)f(x)\mathrm{d}\gamma(x)$
    for every bounded measurable function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$. Most common couplings are independent couplings $\alpha=\mu_{0}\times\mu_{1}$ and so-called ”optimal” transport couplings with respect to the Wasserstein distance. The optimal couplings stand out by the fact that the induced curves are geodesics in the Wasserstein space, which in particular are minimal length curves connecting $\mu_{0}$ and $\mu_{1}$. In an alternative formulation, we may start with random variables with joint law $\alpha=P_{X_{0},X_{1}}$ so that $\mu_{0}=P_{X_{0}}$ and $\mu_{1}=P_{X_{1}}$. Then the induced curve by $\alpha$ becomes $\mu_{t}=P_{X_{t}}$, where
@@
    $\mathcal{K}_{t}(y,\cdot)=e_{t,\sharp}(\alpha^{y}\times\delta_{y})\quad\text{and}\quad v_{t}^{y}=\frac{y-x}{1-t},\quad t\in(0,1)$
@@
Section 7 deals with the actual flow matching, i.e., how we can learn the above vector fields. Actually, we will see that the vector field $v$ of the curve-velocity pair $(\nu_{t},v_{t})$ in the Markov kernel approach appears as minimizer of a functional
@@
which suggests to use the loss function
$\text{FM}_{\alpha_{t}}(\theta)\coloneqq\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|u_{t}^{\theta}(x)-v^{y}_{t}(x)\|^{2}\,\mathrm{d}\alpha_{t}(x,y)\mathrm{d}t$
for determining the parameters $\theta$ of a neural network that approximates $v$. In the original paper [27], the notation ”conditional flow matching” (CFM) was used instead of FM which refers to a ”conditional flow path”. We prefer FM, since we will later deal with Bayesian inverse problems, where the notation of being ”conditional” is occupied in a different way. For the numerical simulation, we have i) to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and ii) to compute $v_{t}^{y}(x)$. Samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are usually available and it should be easy to sample from the distributions $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$. In the special case, where $(\nu_{t},v_{t})=(\mu_{t},v_{t})$ is induced by a coupling $\alpha$, we have a simple formula for $v_{t}^{y}$ for which the loss function becomes
@@ 2.1 Special Metric Spaces
$\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu_{n}=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu\quad\text{for all}\quad\varphi\in C_{0}(\mathbb{R}^{d}).$
@@
$\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu_{n}=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu\quad\text{for all}\quad\varphi\in C_{b}(\mathbb{R}^{d}).$
If $\mu_{n}\in\mathcal{P}(\mathbb{R}^{d})$ with $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, then also $\mu\in\mathcal{P}(\mathbb{R}^{d})$. However, note that $\mathcal{M}(\mathbb{R}^{d})$ is not the dual space of $C_{b}(\mathbb{R}^{d})$, which is indeed the space $\text{rba}(\mathbb{R}^{d})$ of finitely additive set functions. For more information, see [32, Section 4.4].
@@
$\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu=\int\varphi\,\mathrm{d}\nu\text{ for all }\varphi\in C_{0}(\mathbb{R}^{d}).$ (5)
@@ 2.2 Push-Forward Operator
Then, a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is integrable with respect to $T_{\#}\mu$ if $f\circ T$ is integrable with respect to $\mu$ and
$\int_{\mathbb{R}^{d}}f(y)\,\mathrm{d}T_{\#}\mu(y)=\int_{T^{-1}(\mathbb{R}^{d})}f\left(T(x)\right)\,\mathrm{d}\mu(x).$ (6)
@@ 2.4 Couplings and Wasserstein Distance
Indeed, the above minimum is attained, see, e.g. [34, Theorem 1.7], and the set of optimal plans is denoted by $\Gamma_{o}(\mu,\nu)$. The space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a complete metric space. Speaking about $\mathcal{P}_{2}(\mathbb{R}^{d})$, we will always equip it with the $W_{2}$ metric. Then we have for a sequence $(\mu_{n})_{n}$ of measures in $\mathcal{P}_{2}(\mathbb{R}^{d})$ that
@@ 3 Absolutely Continuous Curves in $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$
Recall that in a complete metric space $(X,\text{d})$, a curve $\gamma:I\to X$ is called absolutely continuous, if there exists as $m\in L^{1}([a,b])$ such that
$\text{d}(\gamma(s),\gamma(t))\leq\int_{s}^{t}m(r)\,\mathrm{d}r\quad\text{for all}\quad a\leq s\leq t\leq b.$ (20)
Here $L^{1}([a,b])$ denotes the space of (equivalence classes of) absolutely integrable real-valued function defined on an interval $I$. The metric derivative of $\gamma$ (if it exists) is defined by
$|\gamma^{\prime}|\coloneqq\lim_{h\to 0}\frac{\text{d}\left(\gamma(t),\gamma(t+h)\right)}{|h|}.$
@@
$\int_{I}\left(\sup_{x\in B}\|v_{t}(x)\|+\mathrm{Lip}(v_{t},B)\right)\mathrm{d}t<\infty.$ (28)
@@ 4 Curves Induced by Couplings
Figure 4: Curve induced by $\alpha=(\operatorname{Id},T)_{\sharp}\mu_{0}$ from $\mu_{0}=\delta_{x_{0}}$, resp. $\mu_{0}=\tfrac{1}{2}(\delta_{x_{0}}+\delta_{x_{1}})$ to $\mu_{1}=\tfrac{1}{2}(\delta_{y_{0}}+\delta_{y_{1}})$. In (c), at the crossing time $s$ of the path, there does not exist a map $T_{s}$ that induces an element in $\Gamma_{o}(\mu_{s},\mu_{1})$. Red arrows: vector fields computed via 23
@@
Let $\{g_{n}\}_{n\in\mathbb{N}}\subset C_{b}(\mathbb{R}^{d})$ be a dense subset. Then we obtain for any $h\in C_{b}([0,1]$ and all $n\in\mathbb{N}$ that
@@
The above minimality property, i.e. laying in the tangent space, is quite special for optimal plans. Indeed, by the following example, this is not the case for arbitrary induced velocity fields, see also [28, Example 3.5].
@@
We have already seen that vector fields $v_{t}$ associated to optimal plans are minimal ones, meaning that $v_{t}\in\mathcal{T}_{\mu_{t}}$. This is in general not true for an independent coupling $\alpha=\mu_{0}\times\mu_{1}$ with an arbitrary $\mu_{0}$, see Figure 6. Fortunately, the independent coupling with a Gaussian marginal $\mu_{0}$ has this minimality property as the following example shows, see also [28, Section 5.1].
@@ 2. Concerning relation (62), we verify
$\int_{\mathbb{R}^{d}}g\,v_{t}\,\mathrm{d}\mu_{t}$ $=\int_{\mathbb{R}^{d}}g\,\mathrm{d}e_{t,\sharp}(y-x)\alpha=\frac{1}{(2\pi)^{\frac{d}{2}}}\int g(e_{t}(x,y))(y-x)\text{e}^{-\frac{\|x\|^{2}}{2}}\,\mathrm{d}\mu_{1}(y)\mathrm{d}x$ (70)
@@ 2. For any $\varphi\in C_{c}^{\infty}([0,1]\times\mathbb{R}^{d})$, we obtain
$=\int_{\mathbb{R}^{d}}\int_{0}^{1}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\mathcal{K}_{t}(y,\cdot)(x)\,\mathrm{d}t\,\mathrm{d}\mu_{1}(y)$ (81)
$=\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\alpha_{t}\mathrm{d}t=\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}(\pi^{1}_{\sharp}\alpha_{t})\mathrm{d}t$ (82)
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\nu_{t}(x)\mathrm{d}t,$ (83)

### v4

@@ 1 Introduction
Generative models with neural network approximations have shown impressive results in many applications, in particular in inverse problems and data assimilation. First developments in this direction were generative adversarial networks by Goodfellow et al. [15] in 2014 and variational autoencoders by Kingma and Welling [25] in 2013. In this paper, we are interested in recent flow matching techniques pioneered by Liu et al. [29, 30] in 2022/23 from the point of view of stochastic processes, and by Lipman et al. [28] in 2023 via conditional probability paths. Flow matching is closely related to continuous normalizing flows (CNFs), also called neural ordinary differential equations by Chen et al. [9] in 2018, and to score-based diffusion models introduced by Sohl-Dickstein et al. [38] and Ermon and Song [40] in 2015 and 2019, respectively. The later ones can be considered as an instance of CNFs, but add the perspective of stochastic differential equations. To provide a more complete picture, we will briefly consider these approaches in Sections 10 and 11. Flow matching techniques stand out for their simple applicability and good scaling properties. They can be easily incorporated into other optimization techniques like plug-and-play approaches [31] and can be generalized for solving Bayesian inverse problems as we will see in Section 9. Recently, flow matching was generalized towards generator matching in [20, 22].
@@
Our curves of interest are absolutely continuous curves in the Wasserstein space, which we introduce in Section 3. Due to the special metric of the Wasserstein space, such a curve possesses a vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the curve-velocity pair $(\mu_{t},v_{t})$ fulfills the continuity equation
@@
    Here $v_{t}\mu_{t}=e_{t,\sharp}[(y-x)\alpha]$ is a vector valued measures, where you may note that for $\gamma\in\mathcal{P}_{2}(\mathbb{R}^{d}),f\in L^{1}(\mathbb{R}^{d},\gamma)$ the vector valued measure $f\gamma$ is defined by
    $\int_{\mathbb{R}^{d}}\varphi(x)\mathrm{d}f\gamma(x)=\int_{\mathbb{R}^{d}}\varphi(x)f(x)\mathrm{d}\gamma(x)$
    for every bounded measurable function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$. Most common couplings are independent couplings $\alpha=\mu_{0}\times\mu_{1}$ and so-called ”optimal” transport couplings with respect to the Wasserstein distance. The optimal couplings stand out by the fact that the induced curves are geodesics in the Wasserstein space, which in particular are minimal length curves connecting $\mu_{0}$ and $\mu_{1}$. In an alternative formulation, we may start with random variables with joint law $\alpha=P_{X_{0},X_{1}}$ so that $\mu_{0}=P_{X_{0}}$ and $\mu_{1}=P_{X_{1}}$. Then the induced curve by $\alpha$ becomes $\mu_{t}=P_{X_{t}}$, where
@@
    $\mathcal{K}_{t}(y,\cdot)=e_{t,\sharp}(\alpha^{y}\times\delta_{y})\quad\text{and}\quad v_{t}^{y}=\frac{y-x}{1-t},\quad t\in(0,1)$ (2)
@@
Section 7 deals with the actual flow matching, i.e., how we can learn the above vector field $v_{t}$ by a neural network $u_{t}^{\theta}$. Clearly, e.g. in the Markov kernel approach, the function
$F(\theta)\coloneqq\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|u_{t}^{\theta}(x)-v_{t}(x)\|^{2}\,\mathrm{d}\nu_{t}(x)\mathrm{d}t$
would be a nice loss, but unfortunately we do not have access to $v_{t}$. On the other hand, we know the conditioned velocities $v_{t}^{y}$ from (2). Fortunately, we will see that the vector field $v$ of the curve-velocity pair $(\nu_{t},v_{t})$ appears as minimizer of the functional
@@
or in other words the above loss $F$ coincides up to a constant with the loss function
$\text{FM}_{\alpha_{t}}(\theta)\coloneqq\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|u_{t}^{\theta}(x)-v^{y}_{t}(x)\|^{2}\,\mathrm{d}\alpha_{t}(x,y)\mathrm{d}t.$
In the original paper [28], the notation ”conditional flow matching” (CFM) was used instead of FM which refers to a ”conditional flow path”. We prefer FM, since we will later deal with Bayesian inverse problems, where the notation of being ”conditional” is occupied in a different way. For the numerical simulation, we have on the one hand to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and on the other hand to compute $v_{t}^{y}(x)$. Samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are usually available and it should be easy to sample from the distributions $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$. In the special case, where $(\nu_{t},v_{t})=(\mu_{t},v_{t})$ is induced by a coupling $\alpha$, we have a simple formula for $v_{t}^{y}$ for which the loss function becomes
@@ 2.1 Special Metric Spaces
$\lim_{n\to\infty}\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu_{n}=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu\quad\text{for all}\quad\varphi\in C_{0}(\mathbb{R}^{d}).$
@@
$\lim_{n\to\infty}\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu_{n}=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu\quad\text{for all}\quad\varphi\in C_{b}(\mathbb{R}^{d}).$
If $\mu_{n}\in\mathcal{P}(\mathbb{R}^{d})$ with $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, then also $\mu\in\mathcal{P}(\mathbb{R}^{d})$. However, note that $\mathcal{M}(\mathbb{R}^{d})$ is not the dual space of $C_{b}(\mathbb{R}^{d})$, which is indeed the space $\text{rba}(\mathbb{R}^{d})$ of finitely additive set functions. For more information, see [34, Section 4.4].
@@
$\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\nu\text{ for all }\varphi\in C_{0}(\mathbb{R}^{d}).$ (6)
@@ 2.2 Push-Forward Operator
Then, a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is integrable with respect to $T_{\#}\mu$ if $f\circ T$ is integrable with respect to $\mu$ and
$\int_{\mathbb{R}^{n}}f(y)\,\mathrm{d}T_{\#}\mu(y)=\int_{T^{-1}(\mathbb{R}^{n})}f\left(T(x)\right)\,\mathrm{d}\mu(x).$ (7)
@@ 2.4 Couplings and Wasserstein Distance
Indeed, the above minimum is attained, see, e.g. [36, Theorem 1.7], and the set of optimal plans is denoted by $\Gamma_{o}(\mu,\nu)$. The space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a complete metric space. Speaking about $\mathcal{P}_{2}(\mathbb{R}^{d})$, we will always equip it with the $W_{2}$ metric. Then we have for a sequence $(\mu_{n})_{n}$ of measures in $\mathcal{P}_{2}(\mathbb{R}^{d})$ that
@@ 3 Absolutely Continuous Curves in $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$
Recall that in a complete metric space $(X,\text{d})$, a curve $\gamma:I\to X$ is called absolutely continuous, if there exists as $m\in L^{1}(I)$ such that
$\text{d}(\gamma(s),\gamma(t))\leq\int_{s}^{t}m(r)\,\mathrm{d}r\quad\text{for all}\quad s,t\in I,\;s\leq t.$ (21)
Here $L^{1}([a,b])$ denotes the space of (equivalence classes of) absolutely integrable real-valued function defined on an interval $I$. Then the metric derivative of $\gamma$ defined by
$|\gamma^{\prime}(t)|\coloneqq\lim_{h\to 0}\frac{\text{d}\left(\gamma(t),\gamma(t+h)\right)}{|h|}$
exists a.e. and is the smallest function $m$ in (21). The space of absolutely continuous curves is denoted by $\text{AC}^{1}(X,\mathrm{d})$. If we replace in the above definition $m\in L^{p}(I)$, $1\leq p<\infty$, we obtain the spaces $\text{AC}^{p}(X,\mathrm{d})$.
@@
$\int_{I}\sup_{x\in B}\|v_{t}(x)\|+\mathrm{Lip}(v_{t},B)\,\mathrm{d}t<\infty.$ (29)
@@ 4 Curves Induced by Couplings
Figure 4: Curve induced by $\alpha=(\operatorname{Id},T)_{\sharp}\mu_{0}$ from $\mu_{0}=\delta_{x_{0}}$, resp. $\mu_{0}=\tfrac{1}{2}(\delta_{x_{0}}+\delta_{x_{1}})$ to $\mu_{1}=\tfrac{1}{2}(\delta_{y_{0}}+\delta_{y_{1}})$. In (c), at the crossing time $s$ of the path, there does not exist a map $T_{s}$ that induces an element in $\Gamma_{o}(\mu_{s},\mu_{1})$. Red arrows: vector fields computed via (105)
@@
Let $\{g_{n}\}_{n\in\mathbb{N}}\subset C_{b}(\mathbb{R}^{d})$ be a dense subset. Then we obtain for any $h\in C_{b}([0,1])$ and all $n\in\mathbb{N}$ that
@@
The above minimality property, i.e. laying in the tangent space, is quite special for optimal plans. Indeed, by the following example, this is not the case for arbitrary induced velocity fields, see also [29, Example 3.5].
@@
We have already seen that vector fields $v_{t}$ associated to optimal plans are minimal ones, meaning that $v_{t}\in\mathcal{T}_{\mu_{t}}$. This is in general not true for an independent coupling $\alpha=\mu_{0}\times\mu_{1}$ with an arbitrary $\mu_{0}$, see Figure 6. Fortunately, the independent coupling with a Gaussian marginal $\mu_{0}$ has this minimality property as the following example shows, see also [29, Section 5.1].
@@ 2. Concerning relation (63), we verify
$\int_{\mathbb{R}^{d}}g\,v_{t}\,\mathrm{d}\mu_{t}$ $=\int_{\mathbb{R}^{d}}g\,\mathrm{d}e_{t,\sharp}(y-x)\alpha=\frac{1}{(2\pi)^{\frac{d}{2}}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}g(e_{t}(x,y))(y-x)\text{e}^{-\frac{\|x\|^{2}}{2}}\,\mathrm{d}\mu_{1}(y)\mathrm{d}x$ (71)
@@
Example 4.12.
We consider $\mu_{0}\sim\mathcal{N}(0,1)$, $\mu_{1}\coloneqq\delta_{0}$ and $\alpha=\mu_{0}\times\mu_{1}$. Then, by (62), the curve $\mu_{t}\coloneqq e_{t,\sharp}\alpha$ admits the strictly positive density
$p_{t}(x)=(1-t)^{-1}(2\pi)^{-\frac{1}{2}}{\rm{e}}^{-\frac{x^{2}}{2(1-t)^{2}}},\quad t\in[0,1),~{}x\in\mathbb{R}.$ (75)
Following (63), we calculate
$f_{t}(x)$ $\coloneqq\frac{1-t}{t}\log p_{t}(x)+\frac{1}{2t}x^{2}$ (76)
$=\frac{1-t}{t}\big{(}\log\left((1-t)^{-1}(2\pi)^{-\frac{1}{2}}\right)-\frac{x^{2}}{2(1-t)^{2}}\big{)}+\frac{1}{2t}x^{2}$ (77)
$=\frac{1-t}{t}\log\left((1-t)^{-1}(2\pi)^{-\frac{1}{2}}\right)-\frac{x^{2}}{2t(1-t)}+\frac{1}{2t}x^{2}$ (78)
and the induced velocity field
$v_{t}(x)$ $=\nabla f_{t}=\frac{1}{t}x-\frac{x}{t(1-t)}$ (79)
$=\frac{x}{t-1}.$ (80)
Hence, for any $x\in\mathbb{R}\setminus\{0\}$, the absolute value $|v_{t}(x)|$ explodes for $t\to 1$. Note that despite the exploding velocity field, the velocity of a trajectory remains constant over time. More precisely, the map $T_{t}(x)=(1-t)x$ is the solution of the flow ODE $\partial_{t}\phi(t,x)=v_{t}(\phi(t,x)),\phi(0,x)=x$ and thus the speed of a trajectory starting at $x$ is constant in time and equal to $v_{t}(T_{t}(x))=-x$.
@@ 2. For any $\varphi\in C_{c}^{\infty}([0,1]\times\mathbb{R}^{d})$, we obtain
$=-\int_{\mathbb{R}^{d}}\int_{0}^{1}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\mathcal{K}_{t}(y,\cdot)(x)\,\mathrm{d}t\,\mathrm{d}\mu_{1}(y)$ (88)
$=-\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\alpha_{t}\mathrm{d}t=\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}(\pi^{1}_{\sharp}\alpha_{t})\mathrm{d}t$ (89)
$=-\int_{0}^{1}\int_{\mathbb{R}^{d}}\partial_{t}\varphi(t,x)\,\mathrm{d}\nu_{t}(x)\mathrm{d}t,$ (90)